OpenStudy (anonymous):

For the principal quantum number n=3, how many possible values are there for the angular momentum quantum number (l)? 2 3 6<<< my guess 8

OpenStudy (aaronq):

why 6? what is your reasoning behind it?

OpenStudy (anonymous):

@aaronq nah, its what I'm going with for the moment........

OpenStudy (aaronq):

in other words, what kinds of orbitals can you have when n=3? s, p, d, f?

OpenStudy (anonymous):

you have s and p orbitals, in the n=3 shell? @aaronq

OpenStudy (aaronq):

you can have s, p and d orbitals when n=3. recall that \(0\leq l\leq n-1\). thus when n=3, \(l=2, 1, 0\)

OpenStudy (anonymous):

so 6! @aaronq

OpenStudy (aaronq):

um nope. l = 2, 1, 0 3 values